# NeRF-MAE: Masked AutoEncoders for Self-Supervised 3D Representation Learning for Neural Radiance Fields

### v2

@@ Abstract
Neural fields excel in computer vision and robotics due to their ability to understand the 3D visual world such as inferring semantics, geometry, and dynamics. Given the capabilities of neural fields in densely representing a 3D scene from 2D images, we ask the question: Can we scale their self-supervised pretraining, specifically using masked autoencoders, to generate effective 3D representations from posed RGB images. Owing to the astounding success of extending transformers to novel data modalities, we employ standard 3D Vision Transformers to suit the unique formulation of NeRFs. We leverage NeRF’s volumetric grid as a dense input to the transformer, contrasting it with other 3D representations such as pointclouds where the information density can be uneven, and the representation is irregular. Due to the difficulty of applying masked autoencoders to an implicit representation, such as NeRF, we opt for extracting an explicit representation that canonicalizes scenes across domains by employing the camera trajectory for sampling. Our goal is made possible by masking random patches from NeRF’s radiance and density grid and employing a standard 3D Swin Transformer to reconstruct the masked patches. In doing so, the model can learn the semantic and spatial structure of complete scenes. We pretrain this representation at scale on our proposed curated posed-RGB data, totaling over 1.6 million images. Once pretrained, the encoder is used for effective 3D transfer learning. Our novel self-supervised pretraining for NeRFs, NeRF-MAE, scales remarkably well and improves performance on various challenging 3D tasks. Utilizing unlabeled posed 2D data for pretraining, NeRF-MAE significantly outperforms self-supervised 3D pretraining and NeRF scene understanding baselines on Front3D and ScanNet datasets with an absolute performance improvement of over 20% AP50 and 8% AP25 for 3D object detection.
@@
Radiance Fields 3D Representation Learning Masked Auto-Encoders
Figure 1: Overview: a) We present NeRF-MAE, the first large-scale self-supervised pretraining utilizing Neural Radiance Field’s (NeRF) radiance and density grid as an input modality. Our approach uses a standard 3D Swin Transformer encoder and a voxel decoder to learn a powerful representation in (a) an opacity-aware dense volumetric masked self-supervised learning objective directly in 3D. (b) Our representation, when fine-tuned on a small subset of data, improves many 3D downstream tasks such as 3D object detection, voxel super-resolution, and voxel-labeling.
@@ 1 Introduction
On the other hand, Vision Transformers (ViT) [Dosovitskiy17] have also transformed computer vision in recent years, excelling in pretext task learning [he2021masked, xie2022simmim, baevski2022data2vec]. Their unique approach, encoding visual representations from patches and leveraging self-attention blocks for modelling long-range global information [raghu2021vision, zhai2022scaling, mao2021voxel], sets them apart from conventional CNNs. ViTs, notably when combined with recent advances in pretraining strategies like masked autoencoders (MAE) [he2021masked, zhang2023learning, yu2022point], outperform CNNs in robust feature representation learning, demonstrating state-of-the-art results when fine-tuned for downstream tasks [raghu2021vision, tang2022self, xie2021self].
Because of NeRF’s inherent ability to densely represent a 3D scene from 2D images, this paper explores scaling up their self-supervised pre-training (as we show in Fig. 2 and Sec. 4) through the use of masked autoencoders. The notion of masked auto-encoding is equally suited to NeRF’s radiance and density grid as to images (Fig. 4). As opposed to other 3D representations such as meshes, point clouds, or lidar data which only model surface-level information, are highly irregular data structures and the information density can be extremely uneven [pang2022masked], NeRF’s radiance and density grid is similar in principle to 2D images for masked auto-encoding. This is the case because it provides high information density, spatial data redundancy [he2021masked], and a regular grid ensuring unbiased sampling between occupied and unoccupied areas. Owing to these nice properties, the radiance and density grid obtained from NeRFs have shown to achieve state-of-the-art scene understanding such as segmentation and detection [hu2023nerf, liu2023instance, xu2023nerf].
@@
For training our model, we curate a large-scale dataset for NeRF pre-training, encompassing 4 diverse sources in synthetic and real domains. Our dataset comprises multi-view posed images, corresponding NeRFs, and sampled radiance and density grids from Front3D, ScanNet, Habitat-Matterport3D, and Hypersim, accumulating over 1.6M images and 3500+ scenes (Table. 1 and Figure 3). Our method proves effective, with pre-trained models demonstrating strong generalization across a range of downstream tasks; thus achieving significant improvement over other self-supervised 3D pretraining baselines (§ 4.3) as well as scene-understanding baselines (§ 4.2). Specifically, NeRF-MAE showcases a 21.5% AP50 and 8% AP25 improvement on Front3D and ScanNet 3D Object Bounding Box (OBB) prediction while requiring less than half the data requirement of the state-of-the-art baseline to achieve the same performance.
To summarize, we make the following contributions:
@@
  • large scale pre-training of Neural Radiance Fields on 4 different datasets, over 1.6M images, and 3500+ indoor scenes, using a single model.
@@ 2 Related Works
3D Representation Learning: Supervised learning for 3D data [qi2017pointnet++, irshad2022shapo, yan2022shapeformer, irshad2022centersnap, he2022voxset] has made promising progress. Although powerful, these techniques require expensive 3D labels. Alternatively, self-supervised learning [yan2022implicit, lunayach2023fsd, liu2022masked, jiang2023self, zhang_depth_contrast, weinzaepfel2022croco] and masked autoencoders [he2021masked, tong2022videomae] have emerged as a way to learn strong representations for computer vision tasks without needing labeled data. Follow-up works extended this idea to the 3D domain by directly reconstructing masked 3D coordinates with point clouds [yu2022point, chen2023pimae, pang2022masked, zhang2022point], meshes [liang2022meshmae] or voxels as 3D representations [hess2022masked]. While successful, they require careful adaptation of existing architectures due to irregularity in data structures and uneven information density whereas our approach uses standard Transformer modules, while also requiring readily available and cheaper to obtain posed 2D data compared with lidar scans or accurate mesh reconstructions.
@@
Figure 3: NeRF-MAE dataset mix: a) Multi-view dataset with camera distribution, b) diverse scenes from 4 different sources i.e. Front3D [fu20213d], Hypersim [roberts2021hypersim], HM3D [ramakrishnan2021hm3d] and ScanNet [dai2017scannet] totaling over 3500 scenes and 1.6M images used for pretraining NeRF-MAE using a single model.
@@
NeRF as 3D Scene Representation: Advances in differentiable neural rendering [niemeyer2020differentiable, Chen2019, remelli2020meshsdf, zhang2021holistic, takikawa2021nglod, wang2021neus] have allowed for learning 3D representations using only image supervision. Neural Radiance Fields (NeRFs) [mildenhall2020nerf] is one example that excels in novel-view synthesis. Extensions of NeRF study generalizable scene modelling [irshad2023neo360, yu2020pixelnerf, chen2021mvsnerf, trevithick2021grf, wu2023reconfusion], semantic and open-world 3D understanding [kundu2022panoptic, Siddiqui_2023_CVPR, Zhi:etal:ICCV2021, lerf2023, Engelmann2023openreno], generative modelling [Niemeyer2020GIRAFFE, wang2022clipnerf, instructnerf2023], robotics [zhu2022nice, sucar2021imap, yenchen2021inerf, Ze2023GNFactor, shen2023F3RM], dynamic scenes [pumarola2020d, liu2020neural, yuan2021star, ost2020neural] and compositional representations [ost2020neural, yang2021objectnerf]. Recently, NeRFs have shown competitive performance for supervised 3D tasks such detection [hu2023nerf, xu2023nerf] and segmentation [liu2023instance] when compared with methods that require depth as input [rukhovich2022fcaf3d]. However, little attention has been given to representation learning for NeRFs. Our approach fills this gap by focusing on self-supervised pretraining for neural radiance fields, utilizing NeRF’s dense radiance and density grid for pretraining to improve performance on various downstream 3D tasks.
@@
Vision Transformers: Computer vision has undergone a transformative shift with the emergence of vision transformers [dosovitskiy2020image, raghu2021vision]. Their ability to model global and local contexts has sparked interest in exploring methods to pre-train these backbones. MoCoV3 [chen2021mocov3] and MAE [he2021masked] delved into different aspects of self-supervised Vision Transformers. In particular, MAE [he2021masked] achieved state-of-the-art results by drawing inspiration from BERT [devlin2018bert], which involves randomly masking words in sentences, and using masked image reconstruction. Subsequent research extended this concept to the 3D domain [pang2022masked, tang2022self, liang2022meshmae, hess2022masked], leading to impressive results in downstream 3D vision tasks.
@@ 3 NeRF-MAE: Self-Supervised Pretraining for NeRFs
NeRF-MAE introduces a self-supervised framework aimed at enhancing 3D representation learning in the context of Neural Radiance Fields [mildenhall2020nerf] (NeRFs). Our approach allows learning robust 3D representations from posed RGB images, leading to substantial improvements in various downstream 3D applications, as we show in Sec. 4. As shown in Figures 1 and 5, such a goal is made possible by two major components: (i) An explicit 4D radiance and density grid extraction module in the canonical world frame using a camera-trajectory aware sampling from a fully-trained implicit NeRF model (described in Sec. 3.2) and (ii) A masked self-supervised pretraining module, operating directly on the explicit NeRF’s 4D radiance and density grid, to train a standard 3D SwinTransformer [liu2021swin, yang2023swin3d] encoder and a voxel decoder using an opacity-aware masked reconstruction objective in 3D (discussed in Sec. 3.4). Our NeRF pretraining dataset is described in Sec. 3.3 and key preliminaries to understand our pipeline are detailed in Sec. 3.1.
Figure 4: NeRF-MAE comparison: We show detailed comparison to point-based pretraining approaches in § 4.3. We note their limitation as they model surface-level sparse points, where our approach is similar to images in terms of high-information density and regularity of structure; hence making NeRF-MAE a direct extension of image MAE to 3D. We leverage NeRF’s dense volumetric information along rays and introduce an opacity-aware reconstruction loss, enabling us to achieve superior representation learning.
@@ 3.2 Neural Radiance Field Grid Representation:
Radiance and Density Grid Sampling from Instant-NGP NeRF: NeRF-MAE reconstructs input masked NeRF volumes to pretrain standard 3D Swin Transformers. The first step in this process is to uniformly sample a radiance and density feature volume from a trained NeRF model (Figure  5), in our case, Instant-NGP NeRF [muller2022instant]. Querying a pre-trained NeRF for radiance and density information on a regular grid allows 1. extracting an explicit representation that compactly captures the original 3D scene, 2. is invariant to the NeRF formulation used, and 3. opens up the possibility of utilizing existing deep-learning architectures developed for 3D tasks such as 3D RPN (Section 4.1) for detection and segmentation [hu2023nerf, liu2023instance]. To sample radiance and density information, we query each trained NeRF model from all the training cameras in the traceable scene volume and average the resulting output.
@@ 3.3 NeRF Pretraining Datasets
A key component of our approach is employing a single model to pre-train strong 3D representations in a self-supervised manner from NeRFs on a large number of scenes with varying complexity and realism. As shown in Figure 3, we train our representation in a fully self-supervised manner on the following dataset mix:
Front3D [fu20213d]: A large-scale dataset containing 18k rooms with 3D objects. We extend [hu2023nerf] to render 5x as many scenes, totaling 1.2k scenes comprising 425k images for our pretraining. The rendering time of this dataset is 6 days on 8 A100 GPUs.
@@
For all datasets, extensive cleaning of the bounding box as well as semantic annotations was performed based on manual filtering, class filtering, and size filtering (see supplementary). This cleaning was only carried out for downstream task fine-tuning and not carried out for our pretraining. We will release the complete dataset along with code, upon acceptance, for reproducibility.
@@ 3.4 Masked Pretraining NeRFs
3D Transformer Encoder: The goal of the encoder is to encode the input 3D radiance and density grid ($\mathcal{G}$) into meaningful features at multiple resolutions which can be decoded and used for downstream tasks (Section 4.1). We employ the 3D version [yang2023swin3d, hu2023nerf] of the standard Swin Transformer [liu2021swin] which is obtained by replacing all the 2D operations of the SwinTransfomer with their 3D equivalent, as noted by [yang2023swin3d, hu2023nerf]; for example 3D convolutions and 3D patch merging. We build our 3D auto-encoder architecture based on standard Transformers since NeRF’s radiance and density grid is similar to images and language in terms of information density and which can serve as a unified architecture for various 3D scene understanding tasks operating directly on NeRFs.
@@
Masked Radiance and Density Grid Pretraining Objective: The goal of masked voxel-grid pretraining as applied to Neural Radiance Field (NeRF) is to encode semantic and spatial regions of interest into the network. We use mask volume region reconstruction, similar to [devlin2018bert] for language and [he2021masked] for images. We enforce a faithful and accurate reconstruction of masked patches with a custom loss function suited to NeRF’s unique formulation. Specifically, we employ a combination of opacity and photometric radiance reconstruction loss, both enforced at the volumetric level, where loss $L_{recon}=\mathcal{L}_{\text{rad}}+\mathcal{L}_{\alpha}$ is defined as:

### v3

@@ Abstract
Neural fields excel in computer vision and robotics due to their ability to understand the 3D visual world such as inferring semantics, geometry, and dynamics. Given the capabilities of neural fields in densely representing a 3D scene from 2D images, we ask the question: Can we scale their self-supervised pretraining, specifically using masked autoencoders, to generate effective 3D representations from posed RGB images. Owing to the astounding success of extending transformers to novel data modalities, we employ standard 3D Vision Transformers to suit the unique formulation of NeRFs. We leverage NeRF’s volumetric grid as a dense input to the transformer, contrasting it with other 3D representations such as pointclouds where the information density can be uneven, and the representation is irregular. Due to the difficulty of applying masked autoencoders to an implicit representation, such as NeRF, we opt for extracting an explicit representation that canonicalizes scenes across domains by employing the camera trajectory for sampling. Our goal is made possible by masking random patches from NeRF’s radiance and density grid and employing a standard 3D Swin Transformer to reconstruct the masked patches. In doing so, the model can learn the semantic and spatial structure of complete scenes. We pretrain this representation at scale on our proposed curated posed-RGB data, totaling over 1.8 million images. Once pretrained, the encoder is used for effective 3D transfer learning. Our novel self-supervised pretraining for NeRFs, NeRF-MAE, scales remarkably well and improves performance on various challenging 3D tasks. Utilizing unlabeled posed 2D data for pretraining, NeRF-MAE significantly outperforms self-supervised 3D pretraining and NeRF scene understanding baselines on Front3D and ScanNet datasets with an absolute performance improvement of over 20% AP50 and 8% AP25 for 3D object detection.
@@
NeRF; Masked AutoEncoders; Vision Transformers, Self-Supervised Learning; Representation Learning; 3D Object Detection; Semantic Labelling
Figure 1: NeRF-MAE Overview: The first large-scale self-supervised pretraining utilizing NeRF’s radiance and density grid as an input modality. Our approach uses standard Transformers to learn powerful 3D representations in (a) an opacity-aware dense volumetric masked self-supervised learning objective. (b) when fine-tuned on a small subset of data, our representation improves many 3D downstream tasks such as 3D object detection, super-resolution, and voxel-labeling.
@@ 1 Introduction
Furthermore, Vision Transformers (ViT) [Dosovitskiy17] have also transformed computer vision in recent years [he2021masked, xie2022simmim, baevski2022data2vec]. Their unique approach, encoding visual information from patches and leveraging self-attention blocks for modeling long-range global information [raghu2021vision, zhai2022scaling, mao2021voxel], sets them apart from conventional CNNs. ViTs, notably when combined with recent advances in pretraining strategies like masked auto-encoders (MAE) [he2021masked, zhang2023learning, yu2022point], outperform CNNs in robust feature representation learning, demonstrating state-of-the-art results when fine-tuned for downstream tasks [raghu2021vision, tang2022self, xie2021self].
Because of NeRF’s inherent ability to densely represent a 3D scene from 2D images, this paper explores scaling up their self-supervised pretraining (as we show in Fig. 2 and Sec. 4) through the use of masked autoencoders. The notion of masked auto-encoding is equally suited to NeRF’s radiance and density grid as to images (Fig. 4). As opposed to other 3D representations such as meshes, point clouds, or lidar data which only model surface-level information, are highly irregular data structures and the information density can be extremely uneven [pang2022masked], NeRF’s radiance and density grid is similar in principle to 2D images for masked auto-encoding. This is the case because it provides high information density, spatial data redundancy [he2021masked], and a regular grid ensuring unbiased sampling between occupied and unoccupied areas. Owing to these nice properties, the radiance and density grid obtained from NeRFs have shown to achieve state-of-the-art scene understanding such as segmentation and detection [hu2023nerf, liu2023instance, xu2023nerf].
@@
For training and evaluating our model, we curate a large-scale dataset for NeRF pretraining, encompassing 4 diverse sources in synthetic and real domains. Our dataset comprises multi-view posed images, corresponding NeRFs, and sampled radiance and density grids from Front3D, ScanNet, ARKitScenes, and Hypersim, accumulating over 1.8M images and 3600+ scenes (Table. 1 and Figure 3). Our method proves effective, with pretrained models demonstrating strong generalization across a range of downstream tasks. Our approach, NeRF-MAE, achieves significant improvement over other self-supervised 3D pretraining baselines (§ 4.3) as well as scene-understanding baselines (§ 4.2), specifically achieving a 21.5% AP50 and 8% AP25 improvement on Front3D and ScanNet 3D Object Bounding Box (OBB) prediction while requiring less than half the data requirement of the state-of-the-art baseline to achieve the same performance. To summarize, we make the following contributions:
@@
  • large scale pretraining of Neural Radiance Fields with robust evaluation on diverse datasets, over 1.8M images, and 3600+ indoor scenes, using a single model.
@@ 2 Related Works
3D Representation Learning: Supervised learning for 3D data [qi2017pointnet++, irshad2022shapo, yan2022shapeformer, irshad2022centersnap, he2022voxset] has made promising progress. Although powerful, these techniques require expensive 3D labels. Alternatively, self-supervised learning [yan2022implicit, liu2022masked, jiang2023self, zhang_depth_contrast, weinzaepfel2022croco] and masked autoencoders [he2021masked] learn strong representations for various tasks without needing labeled data. Follow-up works extended this idea to the 3D domain by reconstructing masked 3D coordinates with point clouds [yu2022point, chen2023pimae, pang2022masked, zhang2022point], meshes [liang2022meshmae] or voxels [hess2022masked] as 3D representations. While successful, they require careful adaptation of existing architectures due to irregularity in data structures and uneven information density. Contrastingly, NeRF-MAE uses standard Transformers, while requiring readily available and cheaper-to-obtain posed 2D data compared to lidar scans or accurate mesh reconstructions.
@@
Figure 3: NeRF-MAE dataset mix: a) Multi-view dataset with camera distribution, b) diverse scenes from different sources i.e. Front3D [fu20213d], Hypersim [roberts2021hypersim], HM3D [ramakrishnan2021hm3d] and ScanNet [dai2017scannet] totaling over 3600 scenes and 1.8M images used for pertaining and evaluating NeRF-MAE.
@@
NeRF as 3D Scene Representation: Advances in differentiable rendering [niemeyer2020differentiable, Chen2019, remelli2020meshsdf, zhang2021holistic, takikawa2021nglod, wang2021neus] have allowed for learning 3D representations using only image supervision. Neural Radiance Fields (NeRFs) [mildenhall2020nerf] is one example that excels in novel-view synthesis. Extensions of NeRF study generalizable scene modelling [irshad2023neo360, yu2020pixelnerf, chen2021mvsnerf, trevithick2021grf, wu2023reconfusion], semantic and open-world 3D understanding [kundu2022panoptic, Siddiqui_2023_CVPR, Zhi:etal:ICCV2021, lerf2023, Engelmann2023openreno], generative modelling [Niemeyer2020GIRAFFE, wang2022clipnerf, instructnerf2023], robotics [zhu2022nice, sucar2021imap, yenchen2021inerf, Ze2023GNFactor, shen2023F3RM], dynamic scenes [pumarola2020d, liu2020neural, yuan2021star, ost2020neural] and compositional representations [ost2020neural, yang2021objectnerf]. Recently, NeRFs have shown competitive performance for supervised 3D tasks such detection [hu2023nerf, xu2023nerf] and segmentation [liu2023instance] when compared with methods that require depth as input [rukhovich2022fcaf3d]. However, little attention has been given to representation learning for NeRFs. Our approach fills this gap by focusing on self-supervised pretraining for neural radiance fields, utilizing NeRF’s dense radiance and density grid for pretraining to improve performance on various downstream 3D tasks.
@@
Vision Transformers: Computer vision has undergone a transformative shift with the emergence of vision transformers [dosovitskiy2020image, raghu2021vision]. Their ability to model global and local contexts has sparked interest in exploring methods to pretrain these backbones. MoCoV3 [chen2021mocov3] and MAE [he2021masked] delved into different aspects of self-supervised Vision Transformers. In particular, MAE [he2021masked] achieved state-of-the-art results by drawing inspiration from BERT [devlin2018bert], which involves randomly masking words in sentences, and using masked image reconstruction. Subsequent research extended this concept to the 3D domain [pang2022masked, tang2022self, liang2022meshmae, hess2022masked], leading to impressive results in downstream 3D vision tasks.
@@ 3 NeRF-MAE: Self-Supervised Pretraining for NeRFs
NeRF-MAE introduces a self-supervised framework aimed at enhancing 3D representation learning in the context of Neural Radiance Fields [mildenhall2020nerf]. Our approach allows learning robust 3D representations from posed RGB images, leading to substantial improvements in various downstream 3D applications, as we show in Sec. 4. As shown in Figures 1 and 5, such a goal is made possible by two major components: (i) An explicit 4D radiance and density grid extraction module in the canonical world frame using a camera-trajectory aware sampling from a fully-trained implicit NeRF model (described in Sec. 3.2) and (ii) A masked self-supervised pretraining module, operating directly on the explicit NeRF’s 4D radiance and density grid, to train a standard 3D SwinTransformer [liu2021swin, yang2023swin3d] encoder and a voxel decoder using an opacity-aware masked reconstruction objective in 3D (discussed in Sec. 3.4). Our NeRF pretraining dataset is described in Sec. 3.3 and key preliminaries to understand our pipeline are detailed in Sec. 3.1.
Figure 4: NeRF-MAE Comparison: In § 4.3, we compare our method to point-based pretraining approaches, which are limited by their modeling of surface-level sparse points. In contrast, NeRF-MAE resembles images in terms of high information density and structural regularity, hence positioning it as a direct extension of image MAE to 3D. Leveraging NeRF’s dense volumetric information and opacity-aware reconstruction loss, we achieve superior representation learning.
@@ 3.2 Neural Radiance Field Grid Representation:
Radiance and Density Grid Sampling from Instant-NGP NeRF: NeRF-MAE reconstructs input masked NeRF volumes to pretrain standard 3D Swin Transformers. The first step in this process is to uniformly sample a radiance and density feature volume from a trained NeRF model (Figure  5). Querying a pretrained NeRF for radiance and density information on a regular grid allows 1. extracting an explicit representation that compactly captures the original 3D scene, 2. is invariant to the NeRF formulation used, and 3. opens up the possibility of utilizing existing deep-learning architectures developed for 3D tasks such as 3D RPN (Section 4.1) for detection and segmentation [hu2023nerf, liu2023instance]. To sample radiance and density information, we query each trained NeRF model from all the training cameras in the traceable scene volume and average the resulting output.
@@ 3.3 NeRF Pretraining Datasets
A key component of our approach is employing a single model to pretrain strong 3D representations in a self-supervised manner from NeRFs on a large number of scenes with varying complexity and realism. As shown in Figure 3, we train and evaluate our representation in a fully self-supervised manner on the following dataset mix:
Front3D [fu20213d]: A large-scale dataset containing 18k rooms with 3D objects. We extend [hu2023nerf] to render 5x as many scenes, totaling 1.8k scenes comprising 425k images for our pretraining. The rendering time of this dataset is 6 days on 8 A100 GPUs.
@@
For all datasets, extensive cleaning of the bounding box as well as semantic annotations was performed based on manual filtering, class filtering, and size filtering (for further details, please see supplementary).
@@ 3.4 Masked Pretraining NeRFs
3D Transformer Encoder: The goal of the encoder is to encode the input 3D radiance and density grid ($\mathcal{G}$) into meaningful features at multiple resolutions which can be decoded and used for downstream tasks (Section 4.1). We employ the 3D version [yang2023swin3d, hu2023nerf] of the standard Swin Transformer [liu2021swin] which is obtained by replacing all the 2D operations of the SwinTransfomer with their 3D equivalent, as noted by [yang2023swin3d, hu2023nerf]; for example 3D convolutions and 3D patch merging. We build our 3D auto-encoder architecture based on standard Transformers since NeRF’s radiance and density grid is similar to images in terms of information density and which can serve as a unified architecture for various 3D scene understanding tasks operating directly on NeRFs.
@@
Masked Radiance and Density Grid Pretraining Objective: The goal of masked voxel-grid pretraining as applied to Neural Radiance Field (NeRF) is to encode semantic and spatial regions of interest into the network. We use mask volume reconstruction, similar to [he2021masked] for images. We enforce a faithful and accurate reconstruction of masked patches with a custom loss function suited to NeRF’s unique formulation. Specifically, we employ a combination of opacity and photometric radiance reconstruction loss, both enforced at the volumetric level, where loss $L_{recon}=\mathcal{L}_{\text{rad}}+\mathcal{L}_{\alpha}$ is defined as: